### v1

# Exact Gauss-Newton Optimization for Training Deep Neural Networks

Abstract
We present EGN, a stochastic second-order optimization algorithm that combines the generalized Gauss-Newton (GN) Hessian approximation with low-rank linear algebra to compute the descent direction. Leveraging the Duncan-Guttman matrix identity, the parameter update is obtained by factorizing a matrix which has the size of the mini-batch. This is particularly advantageous for large-scale machine learning problems where the dimension of the neural network parameter vector is several orders of magnitude larger than the batch size. Additionally, we show how improvements such as line search, adaptive regularization, and momentum can be seamlessly added to EGN to further accelerate the algorithm. Moreover, under mild assumptions, we prove that our algorithm converges to an $\epsilon$-stationary point at a linear rate. Finally, our numerical experiments demonstrate that EGN consistently exceeds, or at most matches the generalization performance of well-tuned SGD, Adam, and SGN optimizers across various supervised and reinforcement learning tasks.
1 Introduction
Optimization plays a pivotal role in Machine Learning (ML), with gradient-based methods being at the forefront. Stochastic Gradient Descent (SGD) [61], a first-order stochastic optimization algorithm, and its accelerated versions such as momentum-based approaches [51, 67, 44, 16], adaptive learning rates [22, 30, 81], and a combination of the two [36, 21, 80], have been instrumental in numerous ML applications. For example, in Computer Vision (CV) ResNets [29] are trained with SGD, AdaGrad [22] is used for training recommendation systems [50], language models GPT-3 [12] and LLaMA [69] are optimized with Adam [36] and AdamW [43], respectively. Despite their cheap and relatively easy-to-implement updates, first order-methods (FOMs) suffer from several shortcomings. FOMs are sensitive to hyper-parameter selection, and the optimal hyper-parameter set typically does not transfer well across different problems which leads to a costly procedure of hyper-parameter tuning. Also, FOMs are slow to converge in the flat regions of the loss landscape, where the Hessian is ill-conditioned [62].
Second-order methods (SOMs) incorporate the (approximate) curvature information into the update in order to effectively precondition the gradient vector. In contrast to first-order algorithms, SOMs are shown to be robust to the selection of hyper-parameters [77] and to potentially offer faster convergence [2, 6]. So far, the adoption of second-order methods for ML problems has been limited due to the complexity in calculating and storing the Hessian and the computational load of solving the linear system $\mathbf{H}\mathbf{d}=-\mathbf{g}$, where $\mathbf{H}$ is the (approximate) Hessian matrix, $\mathbf{g}$ is the gradient of the loss function, and $\mathbf{d}$ is the descent direction. Addressing these computational challenges, most approaches use a combination of Hessian approximation and an efficient algorithmic technique for solving the linear system. Common approximations to the Hessian include diagonal scaling [78, 42], the empirical Fisher matrix [60], the quasi-Newton approach [65, 14, 5], and the Gauss-Newton (GN) approximation [24, 70, 13].
In this work, we follow the Gauss-Newton approach to Hessian approximation. Our innovation is in the use of an efficient exact linear algebra identity—the Duncan-Guttman (DG) formula [23, 26]—to speed up the inversion of the Hessian matrix. Compared to Hessian-free optimization (HFO) [46, 47, 37] and Inexact Gauss-Newton (iGN) [70], Exact Gauss-Newton (EGN) solves the system $\mathbf{H}\mathbf{d}=-\mathbf{g}$ exactly with the same algorithmic complexity burden. The closest approach to EGN, to the best of our knowledge, is the application of the Sherman–Morrison-Woodbury (SMW) formula [60]. As we outline below, by using a different matrix identity we are able to solve for the descent direction in fewer matrix operations than SWM, thus reducing the algorithmic complexity.
Our contributions are as follows.
  • We propose the EGN algorithm, which relies on a regularized Gauss-Newton Hessian matrix and exploits the Duncan-Guttman identity to efficiently solve the linear system.
  • We provide a theoretical analysis of the EGN algorithm and establish an algorithmic upper bound on the complexity of EGN in finding an $\epsilon$-stationary point.
  • We evaluate the performance of EGN on several supervised learning and reinforcement learning tasks using various neural network architectures.
2 Preliminaries
2.1 Notation
[…]
2.2 Problem Formulation
We adopt the Empirical Risk Minimization (ERM) framework [49] and consider the problem of finding the weights $\mathbf{w}\in{\mathbb{R}}^{d}$ of a parametric function $\Phi:{\mathbb{R}}^{m}\times{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}^{c}$, e.g., a neural network, such that it minimizes the empirical risk over a dataset $\mathcal{D}$ consisting of $N$ pairs $(\mathbf{y}_{i},\mathbf{x}_{i})$ where $\mathbf{x}_{i}\in{\mathbb{R}}^{m}$ is a vector of features and $\mathbf{y}_{i}\in{\mathbb{R}}^{c}$, $c\geq 1$ is a target vector. We want to solve the following optimization problem:
$\mathbf{w}^{*}\in\operatorname*{arg\,min}\limits_{\mathbf{w\in}\mathbb{R}^{d}}\mathcal{L}_{N}\left(\mathbf{w}\right)\coloneqq{\mathbb{E}}_{\xi\sim P_{\xi}}[{\mathcal{L}}_{N}(\mathbf{w};\xi)],$ (2.1)
where $\xi$ is a random variable with distribution $P_{\xi}$. The objective function is expressed as a finite-sum of functions ${\mathcal{L}}_{N}(\mathbf{w};\xi_{i})$ over the realization $\mathcal{D}$ of $\xi$, and
$\mathcal{L}_{N}\left(\mathbf{w}\right)\coloneqq\frac{1}{N}\sum_{i=1}^{N}\ell\left(\mathbf{y}_{i},\Phi(\mathbf{x}_{i};\mathbf{w})\right),$ (2.2)
is the empirical risk with $\ell\colon{\mathbb{R}}^{c}\times{\mathbb{R}}^{c}\to{\mathbb{R}}$ – a loss function, involving a single pair $\left(\mathbf{y}_{i},\mathbf{x}_{i}\right)$ only.
We assume all these functions to be twice differentiable with respect to $\mathbf{w}$. Note that, while this assumption could be partially relaxed, we stick to it for the sake of simplicity.
2.3 Gradient-based Optimization
Problem (2.1) is typically solved by variations of the Stochastic Gradient Descent (SGD) method by sampling mini-batches $\mathcal{B}_{t}$ from $\mathcal{D}$ rather than processing the entire dataset in each iteration. We denote the loss on a mini-batch $\mathcal{L}_{b}$ as
$\mathcal{L}_{b}\left(\mathbf{w}\right)\coloneqq\frac{1}{b}\sum_{i=1}^{b}\ell\left(\mathbf{y}_{i},\Phi(\mathbf{x}_{i};\mathbf{w})\right),$ (2.3)
where $b$ is the batch size. The iterations take the form
$\mathbf{w}_{t+1}\leftarrow\mathbf{w}_{t}+\alpha_{t}\mathbf{d}_{t},$ (2.4)
where $\alpha_{t}>0$ is a learning rate and $\mathbf{d}_{t}$ is a descent direction obtained from the gradient. In general, we have $\mathbf{d}_{t}=-\mathbf{C}_{t}\mathbf{g}_{t}$ where $\mathbf{C}_{t}$ is a preconditioning matrix that scales, rotates, and shears the gradient of the mini-batch loss $\mathbf{g}\coloneqq\nabla_{\mathbf{w}}\mathcal{L}_{b}$. Notice that by setting $b=1$, $\mathbf{C}_{t}=\mathbf{I}_{d}$ and $\mathbf{d}_{t}=-\mathbf{g}_{t}$ we recover the incremental SGD update [61]. In practice, the preferred training algorithm is often an accelerated version of mini-batch SGD.
Minimizing the quadratic approximation of the batch loss leads to the following linear system
$\mathbf{H}_{t}\mathbf{d}_{t}=-\mathbf{g}_{t},$ (2.5)
where $\mathbf{H}_{t}\coloneqq\nabla_{\mathbf{w}}^{2}\mathcal{L}_{b}$ is the Hessian matrix of the mini-batch loss. Finding the direction $\mathbf{d}_{t}$ by solving the system (2.5) using $\mathbf{H}_{t}$ or its approximation defines the broad spectrum of second-order methods. Setting $\mathbf{C}_{t}=\mathbf{H}^{-1}_{t}$ results in Newton’s method. This method suffers from several drawbacks: (a) one needs to compute second-order derivatives with respect to $\mathbf{w}$; (b) $\mathbf{C}_{t}$ has to be positive-definite to ensure descent; (c) the linear system (2.5) must be solved, which in general scales cubically with the dimension of $\mathbf{w}$; and (d) since $\mathbf{H}_{t}$ is a noisy estimate of the true Hessian of the empirical risk $\mathcal{L}$, $\mathbf{H}^{-1}_{t}$ can result in suboptimal conditioning. These issues and the fact that $\mathbf{w}$ is of rather large dimension have made the direct application of Newton’s method practically irrelevant in ML applications. Indeed, for problems like Large Language Model (LLM) pre-training [71, 12, 69] or CV tasks [29], $\mathbf{w}$ can be extremely high-dimensional, e.g., $314\cdot 10^{9}$ parameters for Grok-1 [76] model, $7\cdot 10^{9}$ to $65\cdot 10^{9}$ parameters for LLaMA-family models [69] and $0.27\cdot 10^{6}$ to $19.4\cdot 10^{6}$ parameters for ResNets [29], which explains why accelerated first-order methods are usually preferred.
In order to make SOMs scalable for ML applications one could instead approximate the inverse Hessian, i.e., $\mathbf{C}_{t}\approx\mathbf{H}^{-1}_{t}$. Examples of such preconditioning include diagonal scaling (e.g., with Hutchinson method [78, 42]) with the idea of extracting the (approximate) diagonal elements of $\mathbf{H}_{t}$ while neglecting the off-diagonal terms; the quasi-Newton approach [14, 5] that approximates the Hessian using the information from past and current gradients; and the Gauss-Newton method [24, 60] that leverages the Jacobian of the residuals, neglecting second-order cross-derivatives, particularly suitable for loss functions structured as (2.2). In this paper, we use the Gauss-Newton Hessian approximation which is shown to provide a good approximation of the true Hessian for ML applications [62, 56].
2.4 Generalized Gauss-Newton Hessian Approximation
We consider the Generalized Gauss-Newton (GGN) Hessian approximation scheme [64, 8, 56] which is suited for both regression and multi-class classification tasks.
The GGN Hessian approximation is constructed using only first-order sensitivities (see derivation in Appendix B) to obtain
$\mathbf{H}^{\mathrm{GN}}=\frac{1}{b}\mathbf{J}^{\top}\mathbf{Q}\mathbf{J}$ (2.6)
where we vertically stack individual Jacobians $\mathbf{J}_{\Phi_{i}}:=\frac{\partial\Phi(\mathbf{x}_{i};\mathbf{w})}{\partial\mathbf{w}}$ for each sample in the batch $\mathcal{B}$ to form $\mathbf{J}=\left[\begin{array}[]{ccc}\mathbf{J}_{\Phi_{1}}&\dots&\mathbf{J}_{\Phi_{b}}\end{array}\right]^{\top}\in{\mathbb{R}}^{bc\times d}$ and construct a block diagonal matrix $\mathbf{Q}=\text{blkdiag}(\mathbf{Q}_{\ell_{1}},\mathbf{Q}_{\ell_{2}},\ldots,\mathbf{Q}_{\ell_{b}})\in{\mathbb{R}}^{bc\times bc}$, where $\mathbf{Q}_{\ell_{i}}=\frac{\partial^{2}\ell(\mathbf{y}_{i},\Phi(\mathbf{x}_{i};\mathbf{w}))}{\partial\Phi^{2}}\in{\mathbb{R}}^{c\times c}$. As pointed out in [63], approximation (2.6) is justified since the true Hessian is dominated by the term $\mathbf{J}^{\top}\mathbf{Q}\mathbf{J}$. Using the inverse of the GGN Hessian as a preconditioner yields the following direction
$\mathbf{d}_{t}^{\mathrm{GN}}=-\left(\frac{1}{b}\mathbf{J}_{t}^{\top}\mathbf{Q}_{t}\mathbf{J}_{t}\right)^{-1}\mathbf{g}_{t},$ (2.7)
which, by defining $\Delta\mathbf{w}=\mathbf{w}-\mathbf{w}_{t}$, corresponds to the solution of the quadratic program
$\mathbf{d}_{t}^{\mathrm{GN}}=\operatorname*{arg\,min}\limits_{\Delta\mathbf{w}}\,\frac{1}{2}\Delta\mathbf{w}^{\top}\underbrace{\frac{1}{b}\mathbf{J}_{t}^{\top}\mathbf{Q}_{t}\mathbf{J}_{t}}_{\mathbf{H}_{t}^{\mathrm{GN}}}\Delta\mathbf{w}+\mathbf{g}_{t}^{\top}\Delta\mathbf{w}.$ (2.8)
The Gauss-Newton step (2.7) solves issues (a) and, partially, (b), since no second-order derivatives need to be computed and the Hessian approximation is positive semi-definite by construction provided that the loss function $\ell\left(\mathbf{y}_{i},\Phi(\mathbf{x}_{i};\mathbf{w})\right)$ is convex. A positive-definite Hessian approximation can easily be obtained from $\mathbf{H}^{\mathrm{GN}}$, e.g., by adding to it a small constant times the identity matrix. This approach is called Levenberg-Marquardt (LM) [53] and is often used in practice, such that
$\mathbf{H}^{\mathrm{LM}}=\frac{1}{b}\mathbf{J}^{\top}\mathbf{Q}\mathbf{J}+\lambda\mathbf{I}_{d}$ (2.9)
with $\lambda>0$. The regularizer $\lambda\mathbf{I}_{d}$ serves a dual purpose: it ensures that the approximate Hessian matrix is invertible and also it helps to avoid over-fitting [49].
The Gauss-Newton update, however, still potentially suffers from issue (c), i.e., the need to solve a linear system, which can be of cubic complexity, and (d) the noise in the Hessian estimate. Since issue (c) is a centerpiece of our method, and issue (d) is an inherent part of stochastic sampling we defer the discussion on both of them to Section 3.
We conclude by examining the GGN update in two common machine learning tasks—regression and multi-class classification—and how it is derived in each case.
Regression
Regression is a predictive modeling task where the objective is to predict a scalar numeric target. For regression tasks we define the loss function $\ell$ as the mean squared error (MSE)
$\ell\left(\mathbf{y}_{i},\Phi(\mathbf{x}_{i};\mathbf{w})\right)\coloneqq\frac{1}{2}\left(\Phi(\mathbf{x}_{i};\mathbf{w})-\mathbf{y}_{i}\right)^{2}.$ (2.10)
We can show (see Appendix B.2) that the gradient and the GN Hessian for the MSE loss read
$\mathbf{g}=\frac{1}{b}\mathbf{J}^{\top}\mathbf{r},\qquad\mathbf{H}^{\text{GN}}=\frac{1}{b}\mathbf{J}^{\top}\mathbf{J},$ (2.11)
where $\mathbf{J}\in{\mathbb{R}}^{b\times d}$ is a matrix of stacked Jacobians, and $\mathbf{r}\in{\mathbb{R}}^{b}$ is a vector of residuals defined as $\mathbf{r}:=\left[\begin{array}[]{ccc}\Phi(\mathbf{x}_{1};\mathbf{w})-\mathbf{y}_{1}&\dots&\Phi(\mathbf{x}_{b};\mathbf{w})-\mathbf{y}_{b}\end{array}\right]^{\top}$.
Multi-class Classification
The task of multi-class classification is to predict a correct class from $c$ classes given a vector of features $\mathbf{x}_{i}$. For such problems the output of the neural network $\Phi$ is a $c$-dimensional vector of prediction scores (logits) $\mathbf{z}_{i}=\Phi(\mathbf{x}_{i};\mathbf{w})$ and the target vector is a one-hot encoded vector $\mathbf{y}_{i}=e_{k}$, where $e_{k}$ denotes column $k$ of the identity matrix $\mathbf{I}_{c}$ and $k$ is the index of the correct class. We define the loss function as a softmax cross-entropy loss (CE)
$\ell\left(\mathbf{y}_{i},\Phi(\mathbf{x}_{i};\mathbf{w})\right)\coloneqq-\sum_{k=1}^{c}\mathbf{y}_{i,k}\log\left(\sigma\left(\mathbf{z}_{i,k}\right)\right),$ (2.12)
where $\sigma(\mathbf{z}_{i,k})=\frac{e^{\mathbf{z}_{i,k}}}{\sum_{j=1}^{c}e^{\mathbf{z}_{i,j}}}$, $\mathbf{y}_{i,k}$ and $\mathbf{z}_{i,k}$ are $k$-th elements of vectors $\mathbf{y}_{i}$ and $\mathbf{z}_{i}$ respectively.
As shown in Appendix B.3, the gradient and the GN Hessian for the CE loss are
$\mathbf{g}=\frac{1}{b}\mathbf{J}^{\top}\mathbf{r},\qquad\mathbf{H}^{\text{GN}}=\frac{1}{b}\mathbf{J}^{\top}\mathbf{Q}\mathbf{J},$ (2.13)
where $\mathbf{J}\in{\mathbb{R}}^{bc\times d}$ is a matrix of stacked Jacobians, $\mathbf{r}\in{\mathbb{R}}^{bc}$ is a vector of (pseudo-)residuals defined as $\mathbf{r}:=\left[\begin{array}[]{ccc}\left(\sigma\left(\Phi(\mathbf{x}_{1};\mathbf{w})\right)-\mathbf{y}_{1}\right)^{\top}&\dots&\left(\sigma\left(\Phi(\mathbf{x}_{b};\mathbf{w})\right)-\mathbf{y}_{b}\right)^{\top}\end{array}\right]^{\top}$, and $\mathbf{Q}\in{\mathbb{R}}^{bc\times bc}$ is a block diagonal matrix of stacked matrices $\mathbf{Q}_{\ell_{i}}$ that each have $\sigma(\mathbf{z}_{i,k})(1-\sigma(\mathbf{z}_{i,k}))$ across the diagonal and $-\sigma(\mathbf{z}_{i,k})\sigma(\mathbf{z}_{i,l})$ off-diagonal.
3 Algorithm
We are now ready to present the EGN algorithm. First, we will discuss how one can efficiently solve the linear system. Then, we will discuss further enhancements to the basic algorithm.
Next, we address issue (c), i.e., the problem of finding the solution of the symmetric linear system $\mathbf{H}_{t}\mathbf{d}_{t}=-\mathbf{g}_{t}$. Substituting the exact Hessian with the regularized Gauss-Newton Hessian yields
$\left(\frac{1}{b}\mathbf{J}_{t}^{\top}\mathbf{Q}_{t}\mathbf{J}_{t}+\lambda_{t}\mathbf{I}_{d}\right)\mathbf{d}_{t}^{\mathrm{LM}}=-\frac{1}{b}\mathbf{J}_{t}^{\top}\mathbf{r}_{t},$ (3.1)
where for the MSE loss $c=1$ and $\mathbf{Q}_{t}=\mathbf{I}_{b}$. Solving (3.1) for $\mathbf{d}_{t}^{\mathrm{LM}}$ requires one to factorize matrix $\mathbf{H}^{\mathrm{LM}}$, carrying a complexity of $\mathcal{O}\left(d^{3}\right)$. We notice, however, that in practice one often has $d\gg bc$, i.e., the parameter vector is of very high dimension, e.g., $d>10^{6}$. In that case, the GN Hessian matrix (2.6) is low-rank by construction. That allows us to transfer the computationally expensive inversion operation from the high-dimensional $d\times d$ space (which is the original dimension of the Hessian) to the low-dimensional $bc\times bc$ space.
To that end, we utilize the Duncan-Guttman identity, originally proven in [23, 26].
Theorem 3.1
Assuming $\mathbf{A}$ and $\mathbf{D}$ are full-rank matrices, the following identity holds
$\left(\mathbf{A}-\mathbf{B}^{\top}\mathbf{D}^{-1}\mathbf{C}\right)^{-1}\mathbf{B}^{\top}\mathbf{D}^{-1}=\mathbf{A}^{-1}\mathbf{B}^{\top}\left(\mathbf{D}-\mathbf{C}\mathbf{A}^{-1}\mathbf{B}^{\top}\right)^{-1}.$ (3.2)
By observing that Equation (3.1) defining the direction $\mathbf{d}^{\mathrm{LM}}$ has the form of the left-hand side of the identity (3.2) we state the following (see proof in Appendix B.4).
Lemma 3.2
The Levenberg-Marquardt direction $\mathbf{d}$ (Equation (3.1)) for both MSE and CE loss functions can be computed using Algorithm 1.
Algorithm 1 EGN direction function
1:  Input: (pseudo-)residuals $\mathbf{r}$, stacked Jacobians $\mathbf{J}$, regularizer $\lambda$, batch size $b$.
2:  Solve the linear system for $\delta$: $\left(\mathbf{Q}\mathbf{J}\mathbf{J}^{\top}+b\lambda\mathbf{I}_{bc}\right)\delta=\mathbf{r}$ // $\mathcal{O}\left(b^{2}c^{2}d+b^{3}c^{3}\right)$
3:  Calculate direction $\mathbf{d}^{\mathrm{LM}}=-\mathbf{J}^{\top}\delta$ // $\mathcal{O}\left(bcd\right)$
4:  Return $\mathbf{d}^{\mathrm{LM}}$
The key property of Algorithm 1 is that the system (2.5) is solved exactly in contrast to approximate (or inexact) solutions that underpin algorithms such as HFO [46, 37], Newton-SGI [6], LiSSA [2], SGN [24] and iGN [70]. The benefits of having the exact solution are analyzed in [6] with exact Newton methods enjoying faster convergence rates than inexact ones. Our experiments (Section 5) also demonstrate that the exact Gauss-Newton solver (EGN) consistently outperforms the inexact version (SGN) across the majority of the problems. The complexity of Algorithm 1 is dominated by the matrix multiplication $\mathbf{J}\mathbf{J}^{\top}$ that costs $\mathcal{O}\left(b^{2}c^{2}d\right)$ as well as solving the linear system of size $bc\times bc$ with complexity $\mathcal{O}\left(b^{3}c^{3}\right)$. Assuming $d>bc$, the overall complexity of Algorithm 1 is $\mathcal{O}\left(b^{2}c^{2}d\right)$.
An alternative approach to solve Equation (3.1) exactly is the SWM formula employed in [60, 13, 31, 4]. Solving (3.1) with the SWM formula yields $\mathbf{d}_{t}=-\frac{1}{\lambda}\left(\mathbf{I}_{d}-\frac{1}{b}\mathbf{J}_{t}^{\top}\mathbf{Q}_{t}\left(\lambda\mathbf{I}_{bc}+\frac{1}{b}\mathbf{J}_{t}\mathbf{J}_{t}^{\top}\mathbf{Q}_{t}\right)^{-1}\mathbf{J}_{t}\right)\mathbf{g}_{t}$. As with EGN, the dominating term is $\mathbf{J}\mathbf{J}^{\top}$. However, even with efficient ordering of operations, there are at least two $\mathcal{O}\left(b^{2}c^{2}d\right)$ matrix multiplications while EGN requires just one. Another important consideration is the non-regularized case, i.e., when $\lambda=0$. Assuming $\mathbf{Q}\mathbf{J}\mathbf{J}^{\top}$ is invertible, EGN produces a valid solution while SWM is not defined in that case. That makes Algorithm 1 a preferred choice for finding the direction (B.6) exactly.
For completeness, we note here that another option to solve (3.1) exactly is through a QR factorization of $\mathbf{J}^{\top}_{t}$ (see Appendix C). The complexity of such an approach is also $\mathcal{O}\left(b^{2}c^{2}d\right)$, dominated by the economy size QR decomposition of $\mathbf{J}^{\top}$. However, performing a QR decomposition is significantly more expensive than performing matrix-matrix multiplications and, in practice, EGN is preferred.
The CG method is an essential part of, e.g., HFO [46, 37], SGN [24], Newton-CG [6], LM [59], and Distributed Newton’s method with optimal shrinkage [82]. The complexity of CG approximately solving Equation (3.1) is $\mathcal{O}(ld^{2})$ where $l$ is the number of CG iterations [53]. A typical number of CG iterations ranges from $3$ in [37] to $50$ in [15]. Note that, unless the number of classes $c$ is high (which is one of the limitations of EGN addressed in Section 5.3), we have $ld^{2}>b^{2}c^{2}d$, such that EGN solves the system both exactly and faster than CG.
Besides estimating the Hessian-adjusted direction, most state-of-the-art SOMs employ strategies to reduce variance of gradient and Hessian estimates, safeguard against exploding gradients, and dynamically adjust hyper-parameters to training steps. In Appendix C we discuss several enhancements to EGN, including momentum acceleration [36, 37, 78] to mitigate issue (d), i.e., the noise in Hessian estimates; line search [17, 72, 59] to ensure steps are sufficiently short to decrease the loss; and adaptive regularization [37, 59, 60] to obtain faster convergence and simplify hyper-parameter tuning. The pseudocode for EGN is outlined in Algorithm 5 in Appendix C.
4 Convergence Analysis
In this section, we analyze the convergence of EGN (Algorithm 1) in the general non-convex setting. In EGN, we consider the sequence of iterates $\{\mathbf{w}_{t}\}_{t\geq 1}$ where each $\mathbf{w}_{t}$ is computed via (2.4) with $\mathbf{d}_{t}\equiv\mathbf{d}_{t}^{\mathrm{LM}}$. We aim to minimize the function ${\mathcal{L}}_{N}(\mathbf{w})$ using, for each realization $\xi\sim P_{\xi}$, the Hessian estimator $\mathbf{H}(\mathbf{w},\xi)$ and the gradient estimator $\mathbf{g}(\mathbf{w},\xi)$. Then, at each iteration $t$, the mini-batch estimates $\mathbf{g}_{t}$ and $\mathbf{H}_{t}$ of the gradient and Hessian are
$\mathbf{g}_{t}=\frac{1}{b}\sum_{\xi_{i}\in{\mathcal{B}}_{t}}\mathbf{g}(\mathbf{w},\xi_{i})\coloneqq\frac{1}{b}\mathbf{J}_{t}^{\top}\mathbf{r}_{t},\quad\mathbf{H}_{t}=\frac{1}{b}\sum_{\xi_{i}\in{\mathcal{B}}_{t}}\mathbf{H}(\mathbf{w},\xi_{i})\coloneqq\frac{1}{b}\mathbf{J}_{t}^{\top}\mathbf{Q}_{t}\mathbf{J}_{t}.$ (4.1)
In our analysis, we make use of the following assumptions.
Assumption 4.1
The function ${\mathcal{L}}_{N}$ is lower-bounded on its domain, i.e., $-\infty<{\mathcal{L}}_{N}^{*}\coloneqq\inf\limits_{\mathbf{w}\in{\mathbb{R}}^{d}}{\mathcal{L}}_{N}(\mathbf{w})$. In addition, ${\mathcal{L}}_{N}$ is twice differentiable with Lipschitz continuous first-order derivatives, i.e., $\exists L_{1}\in\mathbb{R}$ such that
$\left\lVert\nabla{\mathcal{L}}_{N}(\bar{\mathbf{w}})-\nabla{\mathcal{L}}_{N}(\tilde{\mathbf{w}})\right\rVert\leq L_{1}\left\lVert\bar{\mathbf{w}}-\tilde{\mathbf{w}}\right\rVert,\quad\forall\bar{\mathbf{w}},\tilde{\mathbf{w}}\in{\mathbb{R}}^{d}.$ (4.2)
Assumption 4.2
At any iteration $t$, $\mathbf{g}(\mathbf{w}_{t},\xi)$ is an unbiased estimator of $\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t})$, i.e.,
${\mathbb{E}}_{\xi}[\mathbf{g}(\mathbf{w}_{t},\xi)]=\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t}).$ (4.3)
Moreover, we have
${\mathbb{E}}_{\xi}\left[\left\lVert\mathbf{g}(\mathbf{w}_{t},\xi)-\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t})\right\rVert^{2}\right]\leq\sigma_{g}^{2},$ (4.4)
where $\sigma_{g}>0$ is a variance parameter.
Assumption 4.3
At any iteration $t$, $\underaccent{\bar}{\kappa}\mathbf{I}\leq\mathbf{Q}_{t}\leq\bar{\kappa}\mathbf{I}$ with $\bar{\kappa}\geq\underaccent{\bar}{\kappa}\geq 0$. Moreover, $\exists\bar{\sigma},\underaccent{\bar}{\sigma}$ satisfying $0<\underaccent{\bar}{\sigma}\leq\bar{\sigma}$ such that $\underaccent{\bar}{\sigma}\leq\left\lVert\mathbf{J}_{t}\right\rVert\leq\bar{\sigma}$ for all $t\geq 0$.
Assumption 4.4
At any iteration $t$, and for any random matrix $\mathbf{B}_{t}$ satisfying $\mathbf{B}_{t}\succeq\mu\mathbf{I}$ with $\mu>0$, ${\mathbb{E}}\left[\langle\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t}),\mathbf{B}_{t}\mathbf{g}_{t}\rangle|\mathbf{w}_{t}\right]\geq\mu K\|\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t})\|\|{\mathbb{E}}[\mathbf{g}_{t}|\mathbf{w}_{t}]\|$, where $K=(3(b\lambda_{t}+\bar{\kappa}\bar{\sigma}^{2}))/(5b\lambda_{t})$.
In Lemma 4.1, we prove a descent lemma for EGN under the given conditions (see Appendix B.5 for the proof).
Lemma 4.1
Let $\{\mathbf{w}_{t}\}$ be the sequence of iterates generated by (2.4) with $\mathbf{d}_{t}\equiv\mathbf{d}_{t}^{\mathrm{LM}}$ and let Assumptions 4.14.3 hold. Assume further that $\alpha_{t}\leq(b\lambda_{t}+\underaccent{\bar}{\kappa}\underaccent{\bar}{\sigma}^{2})^{2}/(b^{2}\lambda_{t}L_{1})$ holds. Then,
${\mathbb{E}}[{\mathcal{L}}_{N}(\mathbf{w}_{t+1})|\mathbf{w}_{t}]\leq{\mathcal{L}}_{N}(\mathbf{w}_{t})-\frac{\alpha_{t}}{10\lambda_{t}}\left\lVert\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t})\right\rVert^{2}+\frac{\sigma_{g}^{2}\alpha_{t}}{2b\lambda_{t}}.$ (4.5)
Assumption 4.4 is necessary to prove the result in Lemma 4.1 due to the correlation between $\mathbf{B}_{t}$ and $\mathbf{g}_{t}$. If these quantities were not correlated, then one could exploit Assumptions 4.24.3: by taking the expected value and using (4.3), one could then bound $\mathbf{B}_{t}\succeq\mu\mathbf{I}$ to obtain the desired bound with $K=1$. Clearly, that would yield different constants in our next results, but their nature would remain unaltered. Note that this observation could suggest alternative versions of EGN, in which the correlation is removed by construction at the expense of an increased computational burden. A thorough investigation of such schemes is beyond the scope of this paper.
We next show the iteration complexity estimate of EGN to reach an $\epsilon$-stationary point. We provide the proof in Appendix B.6.
Theorem 4.2
Let the assumptions in Lemma 4.1 hold, and assume $\lambda_{t}\equiv\lambda$ is fixed for all $t$. Let $\bar{C}$ be a positive number such that ${\mathcal{L}}_{N}(\mathbf{w}_{t})\leq\bar{C}$ holds for all $t$. Then, if the mini-batch size $b$ is such that $b\geq\lambda^{-1}((5/2)\bar{\kappa}\bar{\sigma}-\underaccent{\bar}{\kappa}\underaccent{\bar}{\sigma})>0$, the sequence $\{\mathbf{w}_{t}\}_{t=0}^{T}$, with $T\geq 1$, satisfies
$\frac{1}{T}\sum_{t=0}^{T-1}{\mathbb{E}}[\|\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t})\|^{2}]\leq\frac{4b^{2}\lambda^{2}L_{1}}{\bar{\kappa}\bar{\sigma}^{2}(b\lambda+\underaccent{\bar}{\kappa}\underaccent{\bar}{\sigma}^{2})T}(\bar{C}-{\mathcal{L}}_{N}^{*})+\frac{5\sigma_{g}^{2}}{b}.$ (4.6)
Corollary 4.3
Suppose that the assumptions in Theorem 4.2 hold. Then, for any $\epsilon\in(0,1)$, the sequence $\{\mathbf{w}_{t}\}$ satisfies $\frac{1}{T}\sum_{t=0}^{T-1}{\mathbb{E}}[\|\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t})\|^{2}]\leq\epsilon$ after at most $T=\mathcal{O}(1/\epsilon)$ iterations, provided that $\sigma_{g}^{2}=\mathcal{O}(1/T)$.
5 Experiments
[…]
5.1 Supervised Learning
For the regression task, we select three datasets: California Housing [55], Superconductivity [27], and Diamonds [74] with $20640$, $21263$, and $53940$ training samples, respectively. For classification, we use the IMDB Reviews dataset [45] containing $25000$ instances of movie reviews. Across all problems, the model $\Phi$ is a Feedforward Neural Network (FFNN) with three dense layers of $32$, $64$ and $32$ units followed by the ReLU activation function with a total of $4449$ parameters ($d=4449$). The loss function during the training is a least-squares loss (2.10) for regression and softmax cross-entropy loss (2.12) for classification. The datasets are split into training and test sets in the proportion of $90/10$ percent. Numerical features are scaled and categorical features are one-hot encoded. To measure the performance, we plot the evolution of the evaluation metric on unseen data (test set) with respect to wall time (in seconds). The evaluation metric is Root Mean Squared Error (RMSE) for regression and accuracy for classification.
Figure 1: Learning curves on the test set for SGD, Adam and EGN. The shaded area represents ±$\pm 1$ standard deviation around the mean (thick line) for $10$ seeds.
Table 1: Performance after Training Completion (Supervised Learning)
Optimizer California Housing Superconduct Diamonds IMDB Reviews
SGD 0.539 $\pm$ 0.006 13.788 $\pm$ 0.698 1008.936 $\pm$ 54.940 0.809 $\pm$ 0.011
Adam 0.519 $\pm$ 0.009 12.052 $\pm$ 0.381 947.258 $\pm$ 130.079 0.820 $\pm$ 0.008
EGN 0.518 $\pm$ 0.009 11.961 $\pm$ 0.207 840.500 $\pm$ 126.444 0.830 $\pm$ 0.001
SGN 0.522 $\pm$ 0.007 12.121 $\pm$ 0.196 998.688 $\pm$ 61.489 0.830 $\pm$ 0.001
The results are presented in Figure 1 and Table 1. On all but the IMDB Reviews dataset EGN has achieved both faster convergence and lower test set error than any other optimization algorithm. On IMDB Reviews SGN is faster than EGN, however, the two solvers achieve the same accuracy after reaching convergence.
5.2 Reinforcement Learning
We demonstrate the application of EGN to reinforcement learning in two scenarios: continuous action spaces with Linear-Quadratic Regulator (LQR) and discrete action spaces using Deep Q-Network (DQN) [48].
Learning LQR Controllers
Given a discrete time-invariant linear system with continuous states and actions, and a quadratic reward function our task is to learn the optimal value function $v^{*}(s)$ and the optimal policy $\pi^{*}(s)$ such that we maximize the cumulative return. Such problems can be solved in a data-driven fashion with the policy iteration procedure [10] (outlined in Appendix D.2). It is well known that the optimal value function is quadratic and the optimal policy function is linear [28]. Consequently, we define $\Phi$ as a quadratic function of states and actions. We track the norm of the difference between the optimal LQR controller calculated analytically knowing the system matrices and the learned weights of the model.
We select two linear systems from the Compleib set of benchmarks [39]. The first system is a deterministic model of a binary distillation tower (BDT) [18], and the second one represents the linearized vertical plane dynamics of an aircraft (UAV) with noise [34]. The results are displayed in the two leftmost charts of Figure 2 and Table 2. Both EGN and SGN outperform first-order methods by a considerable margin, with EGN enjoying slightly faster convergence in both cases, while SGN achieves a marginally lower error on the stochastic LQR upon reaching convergence.
Reinforcement Learning with DQN
Adopting the problem formulation of [48], we aim to learn the weights of a neural network that represents a $Q$-value function $q(s,a)$ that maps states $s$ and actions $a$ into scores ($Q$-values) for a discrete set of actions. Once training is complete, the optimal policy is formed by calculating the $Q$-value of each action and choosing the highest-scoring action.
We build upon CleanRL [33] framework for running RL experiments, selecting two environments: Acrobot-v1 and Freeway-v1. Acrobot-v1 is an OpenAI gym [11] environment with a $6$-dimensional state vector and a set of $3$ discrete actions where the goal is to swing the free end of the connected joints above a given height in as few steps as possible. Freeway-v1 is a MinAtar [79] environment that emulates the original Freeway Atari game. The state is represented by a $10\times 10$ image and there are $3$ discrete actions available. For Acrobot-v1 the network $\Phi$ is a FFNN with three dense layers of $32$, $64$ and $32$ units followed by the ReLU activation function with a total of $4515$ parameters. For Freeway-v1 we design a compact CNN, comprising of a convolutional layer with sixteen $3\times 3$ filters and ReLU activation, followed by flattening, a $64$-unit dense layer with ReLU, and a final dense layer outputting $Q$-values for all actions ($d=103683$).
Figure 2: Learning curves for SGD, Adam and EGN. The shaded area represents ±$\pm 1$ standard deviation around the mean return (thick line) for $10$ seeds.
Table 2: Performance after Training Completion (Reinforcement Learning)
Optimizer BDT UAV Acrobot-v1 Freeway-v1
SGD $64.058\pm 0.768$ $81.378\pm 0.306$ $-90.962\pm 13.484$ $56.996\pm 7.359$
Adam $0.443\pm 0.297$ $0.157\pm 0.100$ $-98.489\pm 20.827$ $55.967\pm 2.251$
EGN $0$$\pm$ 0.000 $0.043\pm 0.031$81.796 $\pm$ 11.778 58.916 $\pm$ 1.946
SGN $0$$\pm$ 0.000 0.033 $\pm$ 0.019 $-114.052\pm 39.914$ $57.424\pm 3.611$
The cumulative returns from each completed episode are recorded and displayed in the rightmost charts of Figure 2. The results for both Acrobot-v1 and Freeway-v1 show no distinct advantage among the solvers, as they all reach similar episodic returns. We notice, however, that EGN slightly outperforms other optimizers by achieving a higher return level at convergence (see Table 2).
5.3 Limitations
While delivering promising results, EGN also has some limitations, as we discuss below.
Explicit gradients
Unlike first-order methods that rely on the average gradient of the batch loss, Gauss-Newton methods require the full Jacobian matrix, which contains the gradients of each sample. As a result, backpropagation for EGN is more time-consuming than for FOMs. Moreover, this can lead to increased GPU memory usage, especially with high-dimensional parameter vectors.
Large batch sizes
In our experiments we observed that the cost of computing the derivatives and the subsequent cost of computing the step are approximately the same. However, for large batch sizes ($b>128$) and parameter dimensions ($d>10^{6}$), we observed that the computational times increased significantly. We suspect this to be related to hardware limitations, and future research will aim to investigate this aspect in further detail.
Large number of classes for multi-class classification
By incorporating the softmax function into the loss equation (2.12) we introduce coupling between the individual outputs of $\Phi$ in the denominator $\sum_{j=1}^{c}e^{\mathbf{z}_{i,j}}$, which makes the Jacobian $\mathbf{J}$ a dense $bc\times d$ matrix. Speeding up the calculation of $\mathbf{J}$ remains an open research question and constitutes a major obstacle in using Gauss-Newton methods for tasks involving a large number of classes, e.g., LLM pre-training. One possible solution could consist in moving the softmax function directly into the model $\Phi$ as the last layer of the network and then only computing the Jacobian wrt the correct class, resulting in $\mathbf{J}\in\mathbb{R}^{b\times d}$. This approach, however, yields a Hessian approximation that becomes singular close to the solution, which results in numerical instabilities that hinder convergence [25]. Another possibility suggested in [54] consists in replacing the softmax cross-entropy loss with the multi-class hinge loss. Although the computation becomes faster for the hinge loss, empirical evidence shows that the test set accuracy upon training completion is higher for the CE loss. Finally, a promising approach is the Gauss-Newton-Bartlett (GNB) estimator proposed by [42], which replaces the exact LM Hessian with an approximation obtained by sampling the subset of predicted labels.
6 Conclusions and Future Work
We presented the EGN algorithm, a stochastic second-order method that efficiently estimates the descent direction by using a low-rank Gauss-Newton Hessian approximation and leveraging the Duncan-Guttman matrix identity. We demonstrated that EGN solves the system $\mathbf{H}^{\text{LM}}\mathbf{d}=-\mathbf{g}$ exactly with less computational burden than other exact Gauss-Newton methods, as well as inexact methods that rely on conjugate gradient iterates. We also proved that under mild assumptions our algorithm converges to an $\epsilon$-stationary point at a linear rate. Our empirical results show that EGN consistently matches or exceeds the generalization performance of well-tuned SGD, Adam, and SGN optimizers across various supervised and reinforcement learning tasks.
Future work will focus on addressing the shortcomings of EGN in classification problems with a large number of classes. A promising direction is to approximate the Gauss-Newton Hessian matrix to avoid computing the full Jacobian of the network, e.g., using techniques such as the Gauss-Newton-Bartlett estimator [42]. Another direction is to study the performance of EGN on larger datasets and more complex models.
Appendix A Related Work
Our method can be viewed within the broader context of approximate second-order stochastic optimization. Some notable approaches include diagonal scaling [78, 42], Krylov subspace descent [73], Hessian-free optimization [46, 47, 37], quasi-Newton approach [65, 14, 5, 75, 41], Gauss-Newton  [7, 24] and Natural Gradient [3, 38] methods. A detailed overview of second-order optimization methods for large-scale machine learning problems can be found in in [8, 66, 77].
Most closely related to our work are the algorithms inspired by the Gauss-Newton approach. Those are the methods that approximate the Hessian of the loss function by utilizing first-order sensitivities only. In practice, the damped version of the Gauss-Newton direction is often calculated, forming the stochastic Levenberg-Marquardt (SLM) group of algorithms. We can classify SLM methods by three dimensions: (a) by the type of the Jacobian estimation algorithm used; (b) by the matrix inversion algorithm; and (c) by additional adaptive parameters and acceleration techniques. Based of this paradigm we summarize selected SLM algorithms in Table 3.
Table 3: A survey on Gauss-Newton methods for large-scale stochastic optimization
Algorithm Jacobian Estimation Solving the Linear System Additional Improvements
SGN [24] Exact via reverse mode autodiff Approximate with CG -
LM [59] Exact via reverse mode autodiff Approximate with CG Line search, momentum, uphill step acceptance
SGN [70] Exact via reverse mode autodiff Approximate with ADPG -
SGN2 [70] Approximate with SARAH estimators Approximate with ADPG -
NLLS1, NLLSL [13] Rank-1, Rank-L approximation Exact with SWM formula -
SMW-GN [60] Exact via reverse mode autodiff Exact with SWM formula Adaptive regularization
EGN (ours) Any Jacobian estimation algorithm (exact via backpropagation is the default) Exact with DG identity Line search, adaptive regularization, momentum
The Jacobian can either be calculated exactly through the reverse mode of automatic differentiation as proposed by, e.g., [24, 13, 70] or be estimated approximately. Low rank Jacobian estimation is suggested by the NLLS1 and NLLSL algorithms [13] with experimental results showing almost on par performance with the exact Jacobian version of the methods. SGN2 [70] uses SARAH estimators for approximating function values and Jacobians with SGN2 performing better than SGN [70] that assumes the exact Jacobian. In [42] the Gauss-Newton-Bartlett (GNB) estimator is introduced to adapt the GN method to large-scale classification problems. EGN does not mandate a specific computation technique for the Jacobian. In our experiments we rely on backpropagation deferring other methods to further research.
Solving $\mathbf{H}\mathbf{d}=-\mathbf{g}$ naively for a neural network with $d$ parameters has complexity $\mathcal{O}\left(d^{3}\right)$. The procedure becomes practically infeasible even for networks of moderate size, so several alternative approaches have been proposed. Following [6], we distinguish between inexact and exact solutions to the linear system. Inexact methods rely on iterative algorithms to solve the system approximately in as few iterations as possible. Among such algorithms we mention the Conjugate Gradient (CG) method used in Hessian-free optimization [47, 37] as well as in GN methods like SGN [24] and LM [59]; the Accelerated Dual Proximal-Gradient (ADPG) method proposed in [70] for SGN and SGN2 solvers; the Stochastic Gradient Iteration approach (SGI), analysed in [6] as part of the Newton-SGI solver. Exact methods are typically based on linear algebra identities. For example, in [60] and [13] the system is solved exactly with the Sherman-Morrison-Woodbury (SMW) formula. Contrarily, we follow  [1] and derive the EGN update formula using the Duncan-Guttman matrix identity [23, 26].
State-of-the-art implementations of the GN algorithm often include additional improvements to address the issues of stochasticity of the Hessian matrix, computational load of calculating the Jacobian and adaptive hyper-parameters tuning. Just as with the gradient, the stochastic sampling introduces noise in the Hessian which leads to the erroneous descent direction. A common solution to combat noisy estimates is to add temporal averaging (or momentum), eg. with exponential moving averages (EMA). Examples of such approach are AdaHessian [78] and Sophia [42] that keep EMA of a diagonal Hessian, as well as  [37] that incorporates momentum into the HFO framework. The idea of re-using the Hessian estimate from previous iterations is formalized in [20] showing that evaluating the Hessian “lazily” once per $k$ iterations significantly reduces the computational burden while at the same time does not degrade the performance that much. Although second-order methods typically require less tuning [7, 24, 77], the SLM approach still requires setting a learning rate $\alpha$ and the regularization parameter $\lambda$. The line search for $\alpha$, common in the deterministic optimization, is problematic in the stochastic setting due to the high variance of the loss gradient norm [17]. Still, there are promising attempts to incorporate line search into quasi-Newton methods [75], HFO [37] as well as Gauss-Newton [59]. Adaptive regularization in a manner similar to the deterministic Levenberg-Marquardt approach is proposed in [31, 60].
Appendix B Proofs and Derivations
B.1 Generalized Gauss-Newton Hessian Approximation
Show that the generalized Gauss-Newton Hessian approximation scheme for the batch loss (2.3) is $\mathbf{H}^{\mathrm{GN}}=\frac{1}{b}\mathbf{J}^{\top}\mathbf{Q}\mathbf{J}$.
Proof. The derivative of the generic batch loss function (2.3) reads
$\frac{\partial}{\partial\mathbf{w}}\mathcal{L}_{b}=\frac{\partial}{\partial\mathbf{w}}\left[\frac{1}{b}\sum_{i=1}^{b}\ell\left(\mathbf{y}_{i},\Phi(\mathbf{x}_{i};\mathbf{w})\right)\right]$
Employing the chain rule $\frac{\partial\ell}{\partial\mathbf{w}}=\frac{\partial\ell}{\partial\Phi}\frac{\partial\Phi}{\partial\mathbf{w}}$ we have
$\frac{\partial}{\partial\mathbf{w}}\mathcal{L}_{b}=\frac{1}{b}\sum_{i=1}^{b}\frac{\partial\ell\left(\mathbf{y}_{i},\Phi(\mathbf{x}_{i};\mathbf{w})\right)}{\partial\Phi}\frac{\partial\Phi(\mathbf{x}_{i};\mathbf{w})}{\partial\mathbf{w}}=\frac{1}{b}\sum_{i=1}^{b}\mathbf{L}_{i}\mathbf{J}_{\Phi_{i}},$
where $\mathbf{L}_{i}=\frac{\partial\ell\left(\mathbf{y}_{i},\Phi(\mathbf{x}_{i};\mathbf{w})\right)}{\partial\Phi}\in{\mathbb{R}}^{1\times c}$ and $\mathbf{J}_{\Phi_{i}}=\frac{\partial\Phi(\mathbf{x}_{i};\mathbf{w})}{\partial\mathbf{w}}\in{\mathbb{R}}^{c\times d}$. We obtain the Hessian by differentiating the gradient
$\mathbf{H}=\frac{\partial}{\partial\mathbf{w}}\left[\frac{1}{b}\sum_{i=1}^{b}\mathbf{L}_{i}\mathbf{J}_{\Phi_{i}}\right].$
Using the product rule we arrive at
$\mathbf{H}=\frac{1}{b}\sum_{i=1}^{b}\left(\frac{\partial}{\partial\mathbf{w}}\left[\mathbf{L}_{i}\right]\mathbf{J}_{\Phi_{i}}+\sum_{k=1}^{c}\mathbf{L}_{i,k}\frac{\partial}{\partial\mathbf{w}}\left[\mathbf{J}_{\Phi_{i,k}}\right]\right),$
where $\mathbf{L}_{i,k}\in{\mathbb{R}}$ is the $k$-th element of the derivative of the loss with respect to $\Phi$, and $\mathbf{J}_{\Phi_{i,k}}\in{\mathbb{R}}^{1\times d}$ is the $k$-th row the Jacobian $\mathbf{J}_{\Phi_{i}}$. We obtain the Gauss-Newton Hessian approximation by neglecting the second-term of $\mathbf{H}$ [56, 62], such that
$\mathbf{H}^{\mathrm{GN}}=\frac{1}{b}\sum_{i=1}^{b}\frac{\partial}{\partial\mathbf{w}}\left[\mathbf{L}_{i}\right]\mathbf{J}_{\Phi_{i}}.$
Given that
$\frac{\partial}{\partial\mathbf{w}}\left[\mathbf{L}_{i}\right]=\mathbf{J}_{\Phi_{i}}^{\top}\frac{\partial^{2}\ell\left(\mathbf{y}_{i},\Phi(\mathbf{x}_{i};\mathbf{w})\right)}{\partial\Phi^{2}}=\mathbf{J}_{\Phi_{i}}^{\top}\mathbf{Q}_{\ell_{i}},$
where $\mathbf{Q}_{\ell_{i}}=\frac{\partial^{2}\ell(\mathbf{y}_{i},\Phi(\mathbf{x}_{i};\mathbf{w}))}{\partial\Phi^{2}}\in{\mathbb{R}}^{c\times c}$ is the second derivative of the loss with respect to the function’s output, we have
$\mathbf{H}^{\mathrm{GN}}=\frac{1}{b}\sum_{i=1}^{b}\mathbf{J}_{\Phi_{i}}^{\top}\mathbf{Q}_{\ell_{i}}\mathbf{J}_{\Phi_{i}}.$
Or using the compact notation
$\mathbf{H}^{\mathrm{GN}}=\frac{1}{b}\mathbf{J}^{\top}\mathbf{Q}\mathbf{J},$
where we vertically stack individual Jacobians $\mathbf{J}_{\Phi_{i}}$ for each sample in the batch $\mathcal{B}$ to form $\mathbf{J}\in{\mathbb{R}}^{bc\times d}$ and form a block diagonal matrix $\mathbf{Q}=\text{blkdiag}(\mathbf{Q}_{\ell_{1}},\mathbf{Q}_{\ell_{2}},\ldots,\mathbf{Q}_{\ell_{b}})\in{\mathbb{R}}^{bc\times bc}$. $\Box$
B.2 Gauss-Newton Hessian of the MSE Loss Function
Show that the gradient of the batch MSE loss is $\mathbf{g}=\frac{1}{b}\mathbf{J}^{\top}\mathbf{r}$.
Proof. We define the residual vector $\mathbf{r}$ as $\mathbf{r}:=\left[\begin{array}[]{ccc}\Phi(\mathbf{x}_{1};\mathbf{w})-\mathbf{y}_{1}&\dots&\Phi(\mathbf{x}_{b};\mathbf{w})-\mathbf{y}_{b}\end{array}\right]^{\top}$. Recall that the stacked Jacobians of the neural network for are denoted by $\mathbf{J}$, with $\mathbf{J}\in{\mathbb{R}}^{b\times d}$ for the regression task. The batch loss (2.3) for MSE is
$\mathcal{L}_{b}\left(\mathbf{w}\right)=\frac{1}{2b}\sum_{i=1}^{b}\left(\Phi(\mathbf{x}_{i};\mathbf{w})-\mathbf{y}_{i}\right)^{2}=\frac{1}{2b}\mathbf{r}^{\top}\mathbf{r}.$
So that
$\frac{\partial}{\partial\mathbf{w}}\left[\mathcal{L}_{b}\right]=\frac{\partial\mathcal{L}_{b}}{\partial\mathbf{r}}\frac{\partial\mathbf{r}}{\partial\Phi}\frac{\partial\Phi}{\partial\mathbf{w}}=\frac{1}{b}\mathbf{r}^{\top}\mathbf{J}.$
Since we define the gradient to be a column vector
$\mathbf{g}=\left(\frac{\partial\mathcal{L}_{b}}{\partial\mathbf{w}}\right)^{\top}=\left(\frac{1}{b}\mathbf{r}^{\top}\mathbf{J}\right)^{\top}=\frac{1}{b}\mathbf{J}^{\top}\mathbf{r}.$
$\Box$
Show that the Hessian of the batch MSE loss is $\mathbf{H}^{\text{GN}}=\frac{1}{b}\mathbf{J}^{\top}\mathbf{J}$.
Proof. We obtain the Hessian by differentiating the gradient of the loss
$\mathbf{H}=\frac{\partial}{\partial\mathbf{w}}\mathbf{g}=\frac{1}{b}\frac{\partial}{\partial\mathbf{w}}\left(\mathbf{J}^{\top}\mathbf{r}\right).$
Using the product rule of vector calculus
$\mathbf{H}=\frac{1}{b}\left(\mathbf{J}^{\top}\frac{\partial}{\partial\mathbf{w}}\left[\mathbf{r}\right]+\left(\frac{\partial}{\partial\mathbf{w}}\left[\mathbf{J}^{\top}\right]\right)\mathbf{r}\right)=\frac{1}{b}\left(\mathbf{J}^{\top}\mathbf{J}+\sum_{i=1}^{b}\frac{\partial^{2}}{\partial\mathbf{w}^{2}}\left[\Phi(\mathbf{x}_{i};\mathbf{w})\right]\mathbf{r}_{i}\right),$
where $\mathbf{r}_{i}\in{\mathbb{R}}$ is the $i$-th element of $\mathbf{r}$.
Neglecting the second term we obtain the Gauss-Newton approximation of the Hessian
$\mathbf{H}^{\text{GN}}=\frac{1}{b}\mathbf{J}^{\top}\mathbf{J}.$
$\Box$
B.3 Gauss-Newton Hessian of the Multi-class Cross-entropy Loss Function
Show that the gradient of the batch multi-class cross-entropy loss is $\mathbf{g}=\frac{1}{b}\mathbf{J}^{\top}\mathbf{r}$.
Proof. We recall from Appendix B.1 that the partial derivative of the generic loss function is
$\frac{\partial}{\partial\mathbf{w}}\mathcal{L}_{b}=\frac{1}{b}\sum_{i=1}^{b}\frac{\partial\ell\left(\mathbf{y}_{i},\Phi(\mathbf{x}_{i};\mathbf{w})\right)}{\partial\Phi}\mathbf{J}_{\Phi_{i}}.$
Assuming that $\mathbf{J}_{\Phi_{i}}$ is calculated during the backpropagation stage, we examine the term $\frac{\partial\ell\left(\mathbf{y}_{i},\Phi(\mathbf{x}_{i};\mathbf{w})\right)}{\partial\Phi}$. The cross-entropy loss function (2.12) is defined as
$\ell\left(\mathbf{y}_{i},\Phi(\mathbf{x}_{i};\mathbf{w})\right)\coloneqq-\mathbf{y}_{i}^{\top}\log\left(\sigma\left(\Phi(\mathbf{x}_{i};\mathbf{w})\right)\right),$
where $\sigma$ is the softmax function defined as
$\sigma(\mathbf{z}_{i,k})=\frac{e^{\mathbf{z}_{i,k}}}{\sum_{j=1}^{c}e^{\mathbf{z}_{i,j}}},$
with a $c$-dimensional vector of prediction scores (logits) $\mathbf{z}_{i}=\Phi(\mathbf{x}_{i};\mathbf{w})$ and a $c$-dimensional vector of probabilities $\mathbf{p}_{i}=\sigma(\mathbf{z}_{i})$. Applying the chain rule and using the shorthand notation we obtain
$\frac{\partial\ell\left(\mathbf{y}_{i},\Phi(\mathbf{x}_{i};\mathbf{w})\right)}{\partial\Phi}=\frac{\partial\ell}{\partial\mathbf{p}_{i}}\frac{\partial\mathbf{p}_{i}}{\partial\mathbf{z}_{i}}\frac{\partial\mathbf{z}_{i}}{\partial\Phi}.$
Differentiating the loss with respect to the softmax function yields
$\frac{\partial\ell}{\partial\mathbf{p}_{i}}=\left[-\frac{\mathbf{y}_{i,1}}{\mathbf{p}_{i,1}},-\frac{\mathbf{y}_{i,2}}{\mathbf{p}_{i,2}},\ldots,-\frac{\mathbf{y}_{i,c}}{\mathbf{p}_{i,c}}\right].$
The derivative of the $k$-th element of the softmax function wrt the $l$-th input $\mathbf{z}_{i,l}$ is split in two cases
$\frac{\partial\mathbf{p}_{i,k}}{\partial\mathbf{z}_{i,l}}=\begin{cases}\mathbf{p}_{i,k}(1-\mathbf{p}_{i,k}),&\text{if }k=l\\ -\mathbf{p}_{i,k}\mathbf{p}_{i,l},&\text{if }k\neq l.\end{cases}$
So that the derivative $\frac{\partial\mathbf{p}_{i}}{\partial\mathbf{z}_{i}}\in{\mathbb{R}}^{c\times c}$ can be structured as
$\frac{\partial\mathbf{p}_{i}}{\partial\mathbf{z}_{i}}=\text{diag}(\mathbf{p}_{i})-\mathbf{p}_{i}\mathbf{p}_{i}^{\top}.$
Combining $\frac{\partial\ell}{\partial\mathbf{p}_{i}}$ and $\frac{\partial\mathbf{p}_{i}}{\partial\mathbf{z}_{i}}$, we obtain
$\frac{\partial\ell}{\partial\mathbf{z}_{i,k}}=-\sum_{l=1}^{c}\left(\frac{\mathbf{y}_{i,l}}{\mathbf{p}_{i,l}}\right)\begin{cases}\mathbf{p}_{i,k}(1-\mathbf{p}_{i,k}),&\text{if }k=l\\ -\mathbf{p}_{i,k}\mathbf{p}_{i,l},&\text{if }k\neq l,\end{cases}$
Which simplifies to
$\frac{\partial\ell}{\partial\mathbf{z}_{i,k}}=-\mathbf{y}_{i,k}+\mathbf{p}_{i,k}=\mathbf{p}_{i,k}-\mathbf{y}_{i,k}.$
Since $\frac{\partial\mathbf{z}_{i}}{\partial\Phi}$ is the identity, we have
$\frac{\partial\ell\left(\mathbf{y}_{i},\Phi(\mathbf{x}_{i};\mathbf{w})\right)}{\partial\Phi}=\mathbf{p}_{i}-\mathbf{y}_{i}.$
The same way as for the MSE loss, we define (pseudo-)residuals $\mathbf{r}_{i}\in{\mathbb{R}}^{c}$ as a column vector of probabilities minus the targets, i.e., $\left(\mathbf{p}_{i}-\mathbf{y}_{i}\right)^{\top}$. Substituting back into the loss derivative
$\frac{\partial\mathcal{L}_{b}}{\partial\mathbf{w}}=\frac{1}{b}\sum_{i=1}^{b}\mathbf{r}^{\top}_{i}\mathbf{J}_{\Phi_{i}}.$
Since we define the gradient to be a column vector
$\mathbf{g}=\left(\frac{\partial\mathcal{L}_{b}}{\partial\mathbf{w}}\right)^{\top}=\left(\frac{1}{b}\sum_{i=1}^{b}\mathbf{r}^{\top}_{i}\mathbf{J}_{\Phi_{i}}\right)^{\top}=\frac{1}{b}\sum_{i=1}^{b}\mathbf{J}_{\Phi_{i}}^{\top}\mathbf{r}_{i}.$
Or in shorthand notation
$\mathbf{g}=\frac{1}{b}\mathbf{J}^{\top}\mathbf{r}$
where for each sample $i$ in the batch $\mathcal{B}$ we vertically stack Jacobians $\mathbf{J}_{\Phi_{i}}$ as well residuals $\mathbf{r}_{i}$ to form $\mathbf{J}\in{\mathbb{R}}^{bc\times d}$ and $\mathbf{r}\in{\mathbb{R}}^{bc}$. $\Box$
Show that the Hessian of the batch multi-class cross-entropy loss is $\mathbf{H}^{\text{GN}}=\frac{1}{b}\mathbf{J}^{\top}\mathbf{Q}\mathbf{J}$.
Proof. Recall that the GN Hessian for the generalized case (Appendix B.1) is
$\mathbf{H}^{\mathrm{GN}}=\frac{1}{b}\sum_{i=1}^{b}\mathbf{J}_{\Phi_{i}}^{\top}\mathbf{Q}_{\ell_{i}}\mathbf{J}_{\Phi_{i}}.$ (B.1)
The second derivative of the loss with respect to the function’s output $\mathbf{Q}_{\ell_{i}}=\frac{\partial^{2}\ell(\mathbf{y}_{i},\Phi(\mathbf{x}_{i};\mathbf{w}))}{\partial\Phi^{2}}\in{\mathbb{R}}^{c\times c}$ for CE loss is
$\mathbf{Q}_{\ell_{i}}=\frac{\partial}{\partial\Phi}\left[\mathbf{p}_{i}-\mathbf{y}_{i}\right].$
Using the gradient of the softmax function derived earlier, we form a symmetric matrix $\mathbf{Q}_{\ell_{i}}$ such that
$\mathbf{Q}_{\ell_{i}}=\left[\begin{array}[]{cccc}\mathbf{p}_{i1}(1-\mathbf{p}_{i1})&-\mathbf{p}_{i1}\mathbf{p}_{i2}&\ldots&-\mathbf{p}_{i1}\mathbf{p}_{ic}\\ -\mathbf{p}_{i2}\mathbf{p}_{i1}&\mathbf{p}_{i2}(1-\mathbf{p}_{i2})&\ldots&-\mathbf{p}_{i2}\mathbf{p}_{ic}\\ \vdots&\vdots&\ddots&\vdots\\ -\mathbf{p}_{ic}\mathbf{p}_{i1}&-\mathbf{p}_{ic}\mathbf{p}_{i2}&\ldots&\mathbf{p}_{ic}(1-\mathbf{p}_{ic})\end{array}\right],$
Or in compact notation
$\mathbf{H}^{\text{GN}}=\frac{1}{b}\mathbf{J}^{\top}\mathbf{Q}\mathbf{J},$
Where $\mathbf{Q}$ is a block diagonal matrix
$\mathbf{Q}=\left[\begin{array}[]{cccc}\mathbf{Q}_{\ell_{1}}&0&0&0\\ 0&\mathbf{Q}_{\ell_{2}}&0&0\\ 0&0&\ddots&0\\ 0&0&0&\mathbf{Q}_{\ell_{b}}\end{array}\right]$
And $\mathbf{J}\in{\mathbb{R}}^{bc\times d}$ is vertically stacked Jacobians $\mathbf{J}_{\Phi_{i}}$.
$\Box$
B.4 Proof of Lemma 3.2
Proof. Substituting $\mathbf{A}=b\lambda\mathbf{I}_{d}$, $\mathbf{B}^{\top}=-\mathbf{J}^{\top}$, $\mathbf{C}=\mathbf{J}$ and $\mathbf{D}=\mathbf{Q}^{-1}$ into Equation (3.2) we get
$\left(b\lambda\mathbf{I}_{d}+\mathbf{J}^{\top}\mathbf{Q}\mathbf{J}\right)^{-1}\mathbf{J}^{\top}\mathbf{Q}=-\left(b\lambda\mathbf{I}_{d}\right)^{-1}\mathbf{J}^{\top}\left(\mathbf{Q}^{-1}+\mathbf{J}\left(b\lambda\mathbf{I}_{d}\right)^{-1}\mathbf{J}^{\top}\right)^{-1}$ (B.2)
Multiplying both sides by $\mathbf{Q}^{-1}\mathbf{r}$ yields
$\left(b\lambda\mathbf{I}_{d}+\mathbf{J}^{\top}\mathbf{Q}\mathbf{J}\right)^{-1}\mathbf{J}^{\top}\mathbf{r}=-\left(b\lambda\mathbf{I}_{d}\right)^{-1}\mathbf{J}^{\top}\left(\mathbf{Q}^{-1}+\mathbf{J}\left(b\lambda\mathbf{I}_{d}\right)^{-1}\mathbf{J}^{\top}\right)^{-1}\mathbf{Q}^{-1}\mathbf{r},$ (B.3)
where the lhs of (B.3) is now equivalent to explicitly solving for $\mathbf{d}^{\mathrm{LM}}$ in Equation (3.1). Simplifying the rhs of (B.3) results in
$\mathbf{d}^{\mathrm{LM}}=-\mathbf{J}^{\top}\left(b\lambda\mathbf{Q}^{-1}+\mathbf{J}\mathbf{J}^{\top}\right)^{-1}\mathbf{Q}^{-1}\mathbf{r}.$ (B.4)
Applying the inverse of a product property, $\mathbf{B}^{-1}\mathbf{A}^{-1}=\left(\mathbf{A}\mathbf{B}\right)^{-1}$, to the expression $\left(b\lambda\mathbf{Q}^{-1}+\mathbf{J}\mathbf{J}^{\top}\right)^{-1}\mathbf{Q}^{-1}$ we have
$\left(b\lambda\mathbf{Q}^{-1}+\mathbf{J}\mathbf{J}^{\top}\right)^{-1}\mathbf{Q}^{-1}=\left(\mathbf{Q}\left(b\lambda\mathbf{Q}^{-1}+\mathbf{J}\mathbf{J}^{\top}\right)\right)^{-1}=\left(b\lambda\mathbf{I}_{bc}+\mathbf{Q}\mathbf{\mathbf{J}}\mathbf{J}^{\top}\right)^{-1}.$ (B.5)
So that the direction can be calculated as
$\mathbf{d}^{\mathrm{LM}}=-\mathbf{J}^{\top}\left(\mathbf{Q}\mathbf{J}\mathbf{J}^{\top}+b\lambda\mathbf{I}_{bc}\right)^{-1}\mathbf{r},$ (B.6)
which corresponds to the procedure outlined in Algorithm 1. $\Box$
B.5 Proof of Lemma 4.1
Proof. From Assumption 4.1, we have
${\mathcal{L}}_{N}(\mathbf{w}_{t+1})\leq{\mathcal{L}}_{N}(\mathbf{w}_{t})+\langle\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t}),\mathbf{w}_{t+1}-\mathbf{w}_{t}\rangle+\frac{L_{1}}{2}\left\lVert\mathbf{w}_{t+1}-\mathbf{w}_{t}\right\rVert^{2}$
$={\mathcal{L}}_{N}(\mathbf{w}_{t})-\alpha_{t}\left\langle\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t}),\left(\frac{1}{b}\mathbf{J}_{t}^{\top}\mathbf{Q}_{t}\mathbf{J}_{t}+\lambda_{t}\mathbf{I}\right)^{-1}\mathbf{g}_{t}\right\rangle+\frac{\alpha_{t}^{2}L_{1}}{2}\left\lVert\left(\frac{1}{b}\mathbf{J}_{t}^{\top}\mathbf{Q}_{t}\mathbf{J}_{t}+\lambda_{t}\mathbf{I}\right)^{-1}\mathbf{g}_{t}\right\rVert^{2}.$ (B.7)
Next, set $\mathbf{B}_{t}=\left(\frac{1}{b}\mathbf{J}_{t}^{\top}\mathbf{Q}_{t}\mathbf{J}_{t}+\lambda_{t}\mathbf{I}\right)^{-1}$ in Assumption 4.4. From Assumption 4.3, we can find $\mathbf{B}_{t}\succeq\mu\mathbf{I}$ with $\mu\coloneqq b/(b\lambda_{t}+\bar{\kappa}\bar{\sigma}^{2})$. Now, using these and Assumption 4.3 in (B.7), we have
${\mathcal{L}}_{N}(\mathbf{w}_{t+1})$ $\leq{\mathcal{L}}_{N}(\mathbf{w}_{t})-\alpha_{t}\left\langle\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t}),\mathbf{B}_{t}\mathbf{g}_{t}\right\rangle+\frac{b^{2}\alpha_{t}^{2}L_{1}}{2(b\lambda_{t}+\underaccent{\bar}{\kappa}\underaccent{\bar}{\sigma}^{2})^{2}}\|\mathbf{g}_{t}\|^{2}.$ (B.8)
Notice that by Assumption 4.2, we get that ${\mathbb{E}}[\mathbf{g}_{t}|\mathbf{w}_{t}]=\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t})$, and hence the inequality in Assumption 4.4 may be written as ${\mathbb{E}}\left[\langle\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t}),\mathbf{B}_{t}\mathbf{g}_{t}\rangle|\mathbf{w}_{t}\right]\geq\mu K\|\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t})\|^{2}$. Taking conditional expectation on both sides of (B.8) with respect to $\xi$ and using Assumption 4.4, we get
${\mathbb{E}}[{\mathcal{L}}_{N}(\mathbf{w}_{t+1})|\mathbf{w}_{t}]\leq{\mathcal{L}}_{N}(\mathbf{w}_{t})-\frac{3\alpha_{t}}{5\lambda_{t}}\|\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t})\|^{2}+\frac{b^{2}\alpha_{t}^{2}L_{1}}{2(b\lambda_{t}+\underaccent{\bar}{\kappa}\underaccent{\bar}{\sigma}^{2})^{2}}{\mathbb{E}}[\|\mathbf{g}_{t}\|^{2}|\mathbf{w}_{t}].$ (B.9)
We have also, using Assumption 4.2,
${\mathbb{E}}[\|\mathbf{g}_{t}\|^{2}|\mathbf{w}_{t}]$ $={\mathbb{E}}[\|\mathbf{g}_{t}-\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t})+\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t})\|^{2}|\mathbf{w}_{t}]$
$={\mathbb{E}}[\|\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t})\|^{2}|\mathbf{w}_{t}]+2{\mathbb{E}}[\langle\mathbf{g}_{t}-\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t}),\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t})\rangle|\mathbf{w}_{t}]$
$\qquad+{\mathbb{E}}[\|\mathbf{g}_{t}-\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t})\|^{2}|\mathbf{w}_{t}]$
$=\|\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t})\|^{2}+{\mathbb{E}}[\|\mathbf{g}_{t}-\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t})\|^{2}|\mathbf{w}_{t}]$
$\leq\|\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t})\|^{2}+\frac{\sigma_{g}^{2}}{b}.$ (B.10)
Now, using (B.10) in (B.9), we get
${\mathbb{E}}[{\mathcal{L}}_{N}(\mathbf{w}_{t+1})|\mathbf{w}_{t}]$ $\leq{\mathcal{L}}_{N}(\mathbf{w}_{t})-\left(\frac{3\alpha_{t}}{5\lambda_{t}}-\frac{b^{2}\alpha_{t}^{2}L_{1}}{2(b\lambda_{t}+\underaccent{\bar}{\kappa}\underaccent{\bar}{\sigma}^{2})^{2}}\right)\|\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t})\|^{2}+\frac{b\sigma_{g}^{2}\alpha_{t}^{2}L_{1}}{2(b\lambda_{t}+\underaccent{\bar}{\kappa}\underaccent{\bar}{\sigma}^{2})^{2}}.$ (B.11)
Using $\alpha_{t}\leq(b\lambda_{t}+\underaccent{\bar}{\kappa}\underaccent{\bar}{\sigma}^{2})^{2}/(b^{2}\lambda_{t}L_{1})$ in (B.11) gives (4.5). $\Box$
B.6 Proof of Theorem 4.2
Proof. Taking conditional expectation on both sides of (4.5), we have
${\mathbb{E}}[\|\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t})\|^{2}]\leq\frac{10\lambda}{\alpha_{t}}\left({\mathbb{E}}[{\mathcal{L}}_{N}(\mathbf{w}_{t})]-{\mathbb{E}}[{\mathcal{L}}_{N}(\mathbf{w}_{t+1})]\right)+\frac{5\sigma_{g}^{2}}{b}.$ (B.12)
Summing (B.12) over $t=0,1,\ldots,T-1$, and using Assumption 4.1, we get
$\sum_{t=0}^{T-1}{\mathbb{E}}[\|\nabla{\mathcal{L}}_{N}(\mathbf{w}_{t})\|^{2}]$ $\leq 10\lambda\sum_{t=0}^{T-1}\frac{1}{\alpha_{t}}\left({\mathbb{E}}[{\mathcal{L}}_{N}(\mathbf{w}_{t})]-{\mathbb{E}}[{\mathcal{L}}_{N}(\mathbf{w}_{t+1})]\right)+\frac{5\sigma_{g}^{2}T}{b}$
$=\frac{10\lambda}{\alpha_{0}}{\mathcal{L}}_{N}(\mathbf{w}_{0})+10\lambda\sum_{t=1}^{T-1}\left(\frac{1}{\alpha_{t}}-\frac{1}{\alpha_{t-1}}\right){\mathbb{E}}[{\mathcal{L}}_{N}(\mathbf{w}_{t})]-\frac{10\lambda}{\alpha_{T-1}}{\mathbb{E}}[{\mathcal{L}}_{N}(\mathbf{w}_{T+1})]$
$\qquad+\frac{5\sigma_{g}^{2}T}{b}$
$\leq\frac{10\lambda}{\alpha_{0}}\bar{C}+10\lambda\bar{C}\sum_{t=1}^{T-1}\left(\frac{1}{\alpha_{t}}-\frac{1}{\alpha_{t-1}}\right)-\frac{10\lambda}{\alpha_{T-1}}{\mathcal{L}}_{N}^{*}+\frac{5\sigma_{g}^{2}T}{b}$
$=\frac{10\lambda}{\alpha_{T-1}}(\bar{C}-{\mathcal{L}}_{N}^{*})+\frac{5\sigma_{g}^{2}T}{b}$
$\leq\frac{4b^{2}\lambda^{2}L_{1}}{\bar{\kappa}\bar{\sigma}^{2}(b\lambda+\underaccent{\bar}{\kappa}\underaccent{\bar}{\sigma}^{2})}(\bar{C}-{\mathcal{L}}_{N}^{*})+\frac{5\sigma_{g}^{2}T}{b},$ (B.13)
where we have used that $\alpha_{T-1}\geq(5\bar{\kappa}\bar{\sigma}^{2}(b\lambda+\underaccent{\bar}{\kappa}\underaccent{\bar}{\sigma}^{2}))/(2b^{2}\lambda L_{1})$ holds under the condition $b\geq\lambda^{-1}((5/2)\bar{\kappa}\bar{\sigma}-\underaccent{\bar}{\kappa}\underaccent{\bar}{\sigma})>0$. Dividing (B.13) through by $T$ gives the result. $\Box$
Appendix C Algorithms
Algorithm 2 Calculate direction using x2 QR factorization (MSE loss)
1:  Input: (pseudo-)residuals $\mathbf{r}$, stacked Jacobians of the model $\mathbf{J}$, regularizer $\lambda$.
2:  Factorize $\mathbf{J}^{\top}$ with economy sized QR: $\mathbf{Q},\mathbf{R}\leftarrow\text{qr}\left(\mathbf{J}^{\top}\right)$
3:  Factorize: $\mathbf{\tilde{Q}},\mathbf{\tilde{R}}\leftarrow\text{qr}\left(\mathbf{R}\mathbf{R}^{\top}+\lambda\mathbf{I}\right)$
4:  Solve the linear system for $\delta$: $\mathbf{\tilde{R}}\delta=\mathbf{\tilde{Q}^{\top}}\mathbf{R}\mathbf{r}$
5:  Calculate $\mathbf{d}^{\mathrm{LM}}=-\mathbf{Q}\delta$
6:  Return $\mathbf{d}^{\mathrm{LM}}$
Momentum
One aspect of second-order methods not yet addressed is issue (d), specifically, the noise in Hessian estimates. Consider the direction $\mathbf{d}_{t}=-\mathbf{H}^{-1}_{t}\mathbf{g}_{t}$, where both the (approximate) Hessian $\mathbf{H}_{t}$ and the gradient $\mathbf{g}_{t}$ are stochastic estimates of the true derivatives of the empirical risk (2.2). By conditioning the gradient with a noisy Hessian inverse we can introduce additional error in the descent direction. A common technique to stabilize the updates and speed up convergence is temporal averaging (or momentum) that consists in combining past and current estimates. For ML applications the common variants include simple accumulation [22], exponential moving average (EMA) [30], bias-corrected EMA [36, 78] and momentum with an extrapolation step [51].
For diagonal scaling methods, including first-order accelerated algorithms [36, 22, 30] and SOMs [78, 42], the accumulated estimates of both first and second moments are kept separately resulting in $\mathcal{O}\left(d\right)$ space complexity. For Gauss-Newton methods we could alternatively reduce the variance of the Jacobian $\mathbf{J}$, e.g., with SVRG [35], SAGA [19] or SARAH [52], which results in a space complexity of $\mathcal{O}\left(bcd\right)$. Another option is to apply momentum to the descent direction $\mathbf{d}_{t}$, explored in [72, 37], which requires storing a vector of size $d$. Since we never explicitly materialize neither the Hessian nor the gradient (Algorithm 1), we follow the latter approach with bias-corrected EMA by default.
Line Search
Instead of setting $\alpha$ and $\lambda$ manually, we can make these hyper-parameters adapt to the training step to eliminate the costly procedure of hyper-parameter tuning. Line search is a popular technique in deterministic optimization [53] which consists in iteratively searching for a learning rate satisfying some minimum criteria (e.g., Wolfe’s conditions). Providing the same theoretical guarantees becomes challenging in the stochastic setting due to noisy gradients, as outlined in [17]. Despite these difficulties, line search has been successfully applied in practice to both stochastic FOMs [72, 57] and SOMs [59, 75]. In EGN we adopt the strategy suggested by [72], which allows to minimize the time spent in evaluating the loss by incorporating a reset strategy in the beginning of each search (Algorithm 3).
Algorithm 3 Armijo Line Search
1:  Input: direction $\mathbf{d}_{t}$, hyper-parameters $\alpha^{\text{max}}$, $\kappa$, $c^{\text{up}}$, $c^{\text{down}}$.
2:  Initialize $\alpha_{t}\leftarrow\min\left\{\alpha^{\text{max}},\alpha_{t-1}c^{\text{up}}\right\}$
3:  while $\mathcal{L}\left(\mathbf{w}_{t+1}\right)>\mathcal{L}\left(\mathbf{w}_{t}\right)+\kappa\alpha_{t}\nabla\mathcal{L}\left(\mathbf{w}_{t}\right)^{\top}\mathbf{d}_{t}$ do
4:     Update $\alpha_{t}\leftarrow\alpha_{t}c^{\text{down}}$
5:     Update $\mathbf{w}_{t+1}\leftarrow\mathbf{w}_{t}+\alpha_{t}\mathbf{d}_{t}$
6:  end while
7:  Return $\alpha_{t}$
Adaptive regularization
Adaptive regularization techniques for stochastic SOMs are explored in [46, 37, 60, 59] modifying the original Levenberg-Marquardt rule to the stochastic setting. The idea is to track $\rho$, defined as the ratio between the decrease in the actual loss function (2.3) and the decrease in the quadratic model $\mathcal{M}\left(\Delta\mathbf{w}\right)=\mathcal{L}_{b}\left(\mathbf{y}_{t},\Phi(\mathbf{x}_{t};\mathbf{w}_{t})\right)+\mathbf{g}_{t}^{\top}\Delta\mathbf{w}+\frac{1}{2b}\Delta\mathbf{w}^{\top}\mathbf{J}_{t}^{\top}\mathbf{Q}_{t}\mathbf{J}_{t}\Delta\mathbf{w}$, where $\Delta\mathbf{w}=\mathbf{w}-\mathbf{w}_{t}$. This yields
$\rho:=\frac{\mathcal{L}_{b}\left(\mathbf{y}_{t},\Phi(\mathbf{x}_{t};\mathbf{w}_{t+1})\right)-\mathcal{L}_{b}\left(\mathbf{y}_{t},\Phi(\mathbf{x}_{t};\mathbf{w}_{t})\right)}{\mathbf{g}_{t}^{\top}\Delta\mathbf{w}+\frac{1}{2b}\Delta\mathbf{w}^{\top}\mathbf{J}_{t}^{\top}\mathbf{Q}_{t}\mathbf{J}_{t}\Delta\mathbf{w}},$ (C.1)
which measures the accuracy of the quadratic model. In case $\rho$ is small or negative, $\mathcal{M}\left(\Delta\mathbf{w}\right)$ provides inaccurate approximation and the value $\lambda$ is increased. In case $\rho$ is large, $\lambda$ is decreased to give more weight to $\mathcal{M}\left(\Delta\mathbf{w}\right)$ (Algorithm 4). As empirically found by [37], compared to the deterministic LM the increase/decrease coefficients need to be less aggressive to reduce the oscillations of $\lambda_{t}$.
Algorithm 4 Adaptive regularization [37]
1:  Input: batch $\mathcal{B}_{t}$, current weights $\mathbf{w}_{t}$, updated weights $\mathbf{w}_{t+1}$.
2:  Calculate $\rho$ according to (C.1)
3:  if $\rho<0.25$ then
4:     $\lambda_{t+1}\leftarrow 1.01\lambda_{t}$
5:  else if $\rho>0.75$ then
6:     $\lambda_{t+1}\leftarrow 0.99\lambda_{t}$
7:  else
8:     $\lambda_{t+1}\leftarrow\lambda_{t}$
9:  end if
10:  Return $\lambda_{t+1}$
Algorithm 5 incorporates all the improvements discussed above, effectively addressing issues (a)-(d) associated with SOMs and eliminating the need for manual hyper-parameter tuning.
Algorithm 5 EGN
1:  Input: training dataset $\mathcal{D}$, initial weights $\mathbf{w}_{0}$, initial regularizer $\lambda_{0}$, momentum strength $\beta$.
2:  Initialize momentum: $\mathbf{m}_{0}=0$
[…]
4:     Sample a mini-batch $\mathcal{B}_{t}$ from $\mathcal{D}$
5:     Estimate $\mathbf{r}_{t}$ and $\mathbf{J}_{t}$ (e.g., via backpropagation)
6:     Find direction $\mathbf{d}_{t}$ via Algorithm 1
7:     Calculate the momentum term: $\mathbf{m}_{t}\leftarrow\beta\mathbf{m}_{t-1}+(1-\beta)\mathbf{d}_{t}$
8:     Update direction: $\mathbf{d}_{t}\leftarrow\frac{\mathbf{m}_{t}}{1-\beta^{t}}$
9:     Line search for $\alpha_{t}$ via Algorithm 3
10:     Update weights: $\mathbf{w}_{t+1}\leftarrow\mathbf{w}_{t}+\alpha_{t}\mathbf{d}_{t}$
11:     Update $\lambda_{t+1}$ via Algorithm 4
12:  end for
13:  Return $\mathbf{w}_{t}$
Appendix D Experiment Details
D.1 Supervised Learning
California Housing [55], a part of the scikit-learn [58] datasets package, consists of $20640$ samples with $8$ numerical features $\mathbf{x}$ encoding relevant information, e.g., location, median income, etc.; and a real-valued target $\mathbf{y}$ representing the median house value in California as recorded by the 1990 U.S. Census.
Superconductivity [27] is a dataset of HuggingFace Datasets [40] that contains $21263$ instances of $79$ numerical attributes (features $\mathbf{x}$) and critical temperatures (target $\mathbf{y}$) of superconductors.
Diamonds [74] is a TFDS [68] dataset containing $53940$ instances of $9$ physical attributes (both numerical and categorical features $\mathbf{x}$) and prices (target $\mathbf{y}$) of diamonds.
IMDB Reviews [45] is a TFDS [68] dataset that contains $25000$ train samples and $25000$ test samples of movie reviews in a text format. Before passing the samples to the model $\Phi$, we pre-process the raw text data with spaCy [32] en_core_web_lg pipeline that converts a text review into a $300$-dimensional vector of numbers.
The optimal sets of hyper-parameters are presented in Table 4.
Table 4: Optimal Hyper-parameters for Supervised Learning Tasks
Optimizer Learning Rate Regularizer Momentum Line Search #CG Iterations
California Housing
SGD 0.03 - - - -
Adam 0.001 - - - -
EGN 0.4 1.0 0.9 False -
SGN 0.2 1.0 - - 5
Superconduct
SGD 0.0003 - - - -
Adam 0.01 - - - -
EGN 0.05 1.0 0.0 False -
SGN 0.1 1.0 - - 10
Diamonds
SGD 2e-8 - - - -
Adam 0.0005 - - - -
EGN 0.0005 1.0 0.0 False -
SGN 0.001 1.0 - - 5
IMDB Reviews
SGD 0.005 - - - -
Adam 0.01 - - - -
EGN 0.01 1.0 0.0 False -
SGN 0.05 1.0 - - 5
D.2 Learning LQR Controllers
Here, we define the problem of learning an LQR controller more formally. Given a discrete time-invariant linear system with continuous states $\mathcal{S}\in{\mathbb{R}}^{n_{s}}$ and actions $\mathcal{A}\in{\mathbb{R}}^{n_{a}}$ of form $T(s,a)=\mathbf{A}s+\mathbf{B}a+e$ and a reward function $r(s,a)=s^{\top}\mathbf{Q}s+a^{\top}\mathbf{R}a$ our task is to learn the optimal value function $v^{*}(s)$ and the optimal policy $\pi^{*}(s)$ by interacting with $T(s,a)$, where $\mathbf{A}\in{\mathbb{R}}^{n_{s}\times n_{s}}$ and $\mathbf{B}\in{\mathbb{R}}^{n_{a}\times n_{s}}$ are system matrices, $e\sim\mathcal{N}\left(0,\Sigma\right)$ is Gaussian noise, $\mathbf{Q}\in{\mathbb{R}}^{n_{s}\times n_{s}}$ is a negative semi-definite state reward matrix and $\mathbf{R}\in{\mathbb{R}}^{n_{a}\times n_{a}}$ is a negative definite action reward matrix.
It is well-known [28] that the optimal value and policy functions have the form:
$v^{*}(s)$ $=s^{\top}\mathbf{P}s+V_{0},$ $\pi^{*}(s)$ $=\mathbf{K}s,$ (D.1)
where $\mathbf{P}\in{\mathbb{R}}^{n_{s}\times n_{s}}$ is a negative semi-definite matrix, $\mathbf{K}\in{\mathbb{R}}^{n_{a}\times n_{s}}$ is a state feedback matrix, and $V_{0}=\gamma\left(1-\gamma\right)^{-1}\text{Tr}\left(\mathbf{P}\Sigma\right)$.
To learn the optimal controller from data we can utilize Generalized Policy Iteration [10, 9] (Algorithm 6).
Algorithm 6 Generalized Policy Iteration for LQR
1:  Input: initial stabilizing policy $\mathbf{K}_{0}$, initial weights $\mathbf{w}_{0}$, learning rate $\alpha$, discount factor $\gamma$, tolerance $\eta=10^{-8}$, LM regularizer $\epsilon$.
2:  Set policy iteration counter $p=1$
3:  repeat
4:     Given $\mathbf{K}_{p-1}$ estimate the corresponding weights $\mathbf{w}$ through a policy evaluation algorithm, eg. Algorithm 7
5:     Convert weights $\mathbf{w}$ to a matrix $\mathbf{M}$
6:     if $\mathbf{M}_{aa}$ is not positive-definite then
7:        Return Error
8:     end if
9:     Improve policy $\mathbf{K}_{p}=-\mathbf{M}_{aa}^{-1}\mathbf{M}_{as}$
10:     Set $p=p+1$
11:  until $\left\|\mathbf{K}_{p}-\mathbf{K}_{p-1}\right\|<\eta$
12:  Return $\mathbf{K}_{p}$
Algorithm 7 Policy Evaluation for LQR
1:  Input: policy $\mathbf{K}$, initial weights $\mathbf{w}_{0}$, learning rate $\alpha$, discount factor $\gamma$, tolerance $\eta=10^{-8}$.
2:  Set policy evaluation counter $i=1$
3:  Initialize $S_{1}$
4:  repeat
5:     Choose action $A_{i}\sim\pi(S_{i})$ by following an exploratory policy $\pi(s)=\mathbf{K}s+e$
6:     Execute action $A_{i}$ and observe $R_{i},S_{i+1}$
7:     Obtain $A^{\prime}$ by following a greedy policy $\pi(s)=\mathbf{K}s$
8:     Convert $[S_{i},A_{i}]$ and $[S_{i+1},A^{\prime}]$ to quadratic feature vectors $\mathbf{x}$ and $\mathbf{x^{\prime}}$
9:     Calculate $\mathbf{d}_{i}=\left(R_{i}+\mathbf{w}_{i-1}^{\top}(\gamma\mathbf{x}^{\prime}-\mathbf{x})\right)\mathbf{x}$
10:     Update weights $\mathbf{w}_{i}\leftarrow\mathbf{w}_{i-1}+\alpha_{i}\mathbf{d}_{i}$
11:     Set $i=i+1$
12:  until $\left\|\mathbf{w}_{i}-\mathbf{w}_{i-1}\right\|_{\infty}<\eta$
13:  Return $\mathbf{w}_{i}$
BDT
System matrices for the model of a binary distillation tower (BDT) [18] are defined as follows:
$\tiny\mathbf{A}=\left[\begin{array}[]{ccccccccccc}-0.014&0.0043&0&0&0&0&0&0&0&0&0\\ 0.0095&-0.0138&0.0046&0&0&0&0&0&0&0&0.0005\\ 0&0.0095&-0.0141&0.0063&0&0&0&0&0&0&0.0002\\ 0&0&0.0095&-0.0158&0.011&0&0&0&0&0&0\\ 0&0&0&0.0095&-0.0312&0.015&0&0&0&0&0\\ 0&0&0&0&0.0202&-0.0352&0.022&0&0&0&0\\ 0&0&0&0&0&0.0202&-0.0422&0.028&0&0&0\\ 0&0&0&0&0&0&0.0202&-0.0482&0.037&0&0.0002\\ 0&0&0&0&0&0&0&0.0202&-0.0572&0.042&0.0005\\ 0&0&0&0&0&0&0&0&0.0202&-0.0483&0.0005\\ 0.0255&0&0&0&0&0&0&0&0&0.0255&-0.0185\end{array}\right]$
$\tiny\mathbf{B}=\left[\begin{array}[]{ccc}0&0&0\\ 0.000005&-0.00004&0.0025\\ 0.000002&-0.00002&0.005\\ 0.000001&-0.00001&0.005\\ 0&0&0.005\\ 0&0&0.005\\ -0.000005&0.00001&0.005\\ -0.00001&0.00003&0.005\\ -0.00004&0.000005&0.0025\\ -0.00002&0.000002&0.0025\\ 0.00046&0.00046&0\end{array}\right].$
The matrices represent a continuous system. We discretize the system using a Zero-Order Hold (ZOH) method with a sampling rate of $\Delta T=0.1s$.
UAV
System matrices for the linearized vertical plane dynamics of an aircraft (UAV) [34] are defined as follows:
$\mathbf{A}=\left[\begin{array}[]{ccccc}0&0&1.132&0&-1\\ 0&-0.0538&-0.1712&0&0.0705\\ 0&0&0&1&0\\ 0&0.0485&0&-0.8556&-1.013\\ 0&-0.2909&0&1.0532&-0.6859\end{array}\right]\quad\mathbf{B}=\left[\begin{array}[]{ccc}0&0&0\\ -0.12&1&0\\ 0&0&0\\ 4.419&0&-1.665\\ 1.575&0&-0.0732\end{array}\right].$
The matrices represent a continuous system. We discretize the system using a Zero-Order Hold (zoh) method with a sampling rate of $\Delta T=0.1s$.
The optimal sets of hyper-parameters for LQR are presented in Table 5.
Table 5: Optimal Hyper-parameters for BDT and UAV
Optimizer Learning Rate Regularizer Momentum Line Search #CG Iterations
BDT
SGD 0.0000005 - - - -
Adam 0.1 - - - -
EGN 1.0 1.0 0.0 False -
SGN 1.0 1.0 - - 10
UAV
SGD 0.00008 - - - -
Adam 0.02 - - - -
EGN 0.2 1.0 0.0 False -
SGN 1.0 1.0 - - 10
D.3 Reinforcement Learning with DQN
Acrobot
Acrobot-v1 is an OpenAI gym [11] environment where the goal is to swing the free end of the connected joints above a given height in as few steps as possible. Transitions to any non-terminal state yield reward $R_{t}=-1$.
Freeway
Freeway-v1 is a part of MinAtar [79] package that emulates the original Atari Freeway game which play out on a $10\times 10$ grid. The goal is to reach the top of the screen starting at the bottom of the screen maneuvering the obstacles appearing on the screen. A reward of $+1$ is given upon reaching the top of the screen.
The optimal sets of hyper-parameters for reinforcement learning with DQN are presented in Table 6.
Table 6: Optimal Hyper-parameters for Acrobot-v1 and Freeway-v1
Optimizer Learning Rate Regularizer Momentum Line Search #CG Iterations
Acrobot-v1
SGD 0.001 - - - -
Adam 0.0003 - - - -
EGN 0.1 1.0 0.0 False -
SGN 0.005 1.0 - - 3
Freeway-v1
SGD 0.1 - - - -
Adam 0.0003 - - - -
EGN 0.4 1.0 0.0 False -
SGN 0.5 1.0 - - 5